OpenStudy (anonymous):

Mike participated in a bike race. he covered 24 kilometers in 2 hours and the next 18 kilometers in 1 hour. What is the average speed per hour for the total journey? A. 9 kilometers B. 6 kilometers c. 12 kilometers D. 14 kilometers

OpenStudy (anonymous):

average speed = total distance travellled / total time here total distance = 24+ 18 =42 km and total time = 2+1 =3 hr hope u can do the rest..